# Differential Privacy of Cross-Attention with Provable Guarantee

### v1

@@ E.2 High Dimension Single Data Structure
Theorem E.3 (DPTreeHighDim data structure).
@@
    • $z$ is $(\epsilon,\delta+\delta^{\prime})$-DP private of $\sum_{i\in[n]}w_{i}\|y-x_{i}\|_{1}$,
@@
For the runtime analysis, since we loop data structure DPTreeDistance $d$ times, an additional $d$ factor will appear for both initialization and query time complexity. The DP is proved by Lemma E.1. The accuracy is proved by Lemma E.2. ∎
Appendix F Adaptive Query
In this section, we introduce how we can solve the adaptive query problem by our algorithm, using some tools from [75]. Our idea is that, if we can prove that our algorithm can solve any query in the query space with certain error. Then, since adaptive query must lie in this space, we can handle adaptive query. In Section F.1, we show how we can boost the constant probability of our algorithm to high probability. In Section F.2, we show how we can apply the notion of $\epsilon_{0}$-net and bound all query points in net. In Section F.3, we show how we can bound all points in the query space by introducing an additive error. In Section F.4, we examine the effects of different norms on our adaptive query proof.
First, from Theorem E.3, given query $y\in[0,R]^{d},\alpha\in(0,1)$ we have DistanceQuery$(y,\alpha)$ that can solve $d$-dimension weighted $\ell_{1}$-distance problem with constant probability $0.99$. Now we show how to improve it to solve adaptive query problem.
F.1 Boost the Constant Probability to High Probability
We can repeat the data structure multiple times and take the median to boost the constant probability using Chernoff bound from Lemma A.2.
Lemma F.1 (Using Chernoff bound to boost the probability).
@@ F.1 Boost the Constant Probability to High Probability
Then for each fixed query point $y$, we can have the median of $l$ responses is $(\epsilon,\delta+\delta^{\prime})$-DP and the error is upper bounded by
$\alpha A_{*}+B$
with probability $1-p_{f}$.
@@
By basic composition Fact 3.10, we prove the DP. Similar to the proof of Theorem B.2, we prove the error by Chernoff bound (Lemma A.2). ∎
F.2 From Each Fixed Query Point to All On-net Points
@@
Definition F.2 ($\ell_{p}$ $\epsilon_{0}$-net, see Definition 4.2.1 in [85]).
@@
Fact F.3 ($\ell_{\infty}$ $\epsilon_{0}$-net).
@@
Fact F.4 ($\ell_{2}$ $\epsilon_{0}$-net, see Lemma 5 in [88]).
@@
Fact F.5 ($\ell_{1}$ $\epsilon_{0}$-net, see Theorem 2 in [43]).
@@
Lemma F.6 (From for each query point to for all points in net).
@@
Then with probability $1-p_{f}$, for all query points $y\in N$, we can have the median of $l$ responses is $(\epsilon,\delta+\delta^{\prime})$-DP and the error is upper bounded by
$\alpha A_{*}+B.$
@@
By basic composition Fact 3.10, we prove the DP. From Lemma F.1, we know for each $y\in N$, the error is upper bounded by $\alpha A_{*}+B$ with probability $1-p_{f}/|N|$.
@@
F.3 From Net Points to All Points
@@
Lemma F.7 (Lipschitz of query function).
@@
Lemma F.8 (From points in net to all points in query space).
@@
Then with probability $1-p_{f}$, for all query points $q\in{\mathcal{B}}$, there exists a point $y\in N$ which is the closest to $q$, we can have the median of $l$ responses is $(\epsilon,\delta+\delta^{\prime})$-DP and the error satisfy
@@
By basic composition Fact 3.10, we prove the DP.
@@
From Lemma F.1, with $l=O(\log(|N|/p_{f}))$ we know
@@
where the first step follows from definition of $l$, the second step follows from Fact F.3.
We condition on event $E$ to be held. Then, by definition of $\ell_{\infty}$ $\epsilon_{0}$-net (see Definition F.2), for each $q\notin N$, there exists $y\in N$ such that
@@
where the first step follows from Lemma F.7, the second step follows from $\|x\|_{1}\leq d\|x\|_{\infty}$ for $x\in\mathbb{R}^{d}$, and the last step follows from Eq. (3).
@@
where the first step follows from triangular inequality, the second step follows from Eq. (F.3), the third step follows from Lemma F.6, and the last step follows from Eq. (F.3).
@@
F.4 Effect of Different Norms on the Result
In the above proof, we have two different measure spaces, i.e. $\ell_{\infty}$ distance of $\epsilon_{0}$-net (Definition F.2) and $\ell_{1}$ Lipschitz (Lemma F.7).
@@ F.4 Effect of Different Norms on the Result
For different norms, the only differences in the proofs will be Lipschitz smoothness in Eq. (F.3) and the cardinality of $\epsilon_{0}$-net, i.e. $|N|$ in Fact F.3.
Lemma F.9.
If we use $\ell_{\infty}$ $\epsilon_{0}$-net and use $\ell_{1}$ Lipschitz in Lemma F.8, we have copies of data structure
$l=O(d\log(nR/p_{f}))$
@@
If we use $\ell_{\infty}$ to bound the distance to net, Eq. (F.3) is:
@@
where the first step follows from Lemma F.7, the second step follows from $\|x\|_{1}\leq d\|x\|_{\infty}$ for $x\in\mathbb{R}^{d}$, and the last step follows from $\ell_{\infty}$ $\epsilon_{0}$-net.
Then, Eq. (F.3) is
@@
For $\ell_{\infty}$ distance, we have $|N|\leq(5R/\epsilon_{0})^{d}$ in Fact F.3.
We can choose $\epsilon_{0}=\Theta(1/n)$ to hide $nR_{w}\cdot d\cdot\epsilon_{0}$ term in $B$ in Lemma F.8. Thus,
@@
Lemma F.10.
If we use $\ell_{2}$ $\epsilon_{0}$-net and use $\ell_{1}$ Lipschitz in Lemma F.8, we have copies of data structure
$l=O(d\log(nR/p_{f}))$
@@
If we use $\ell_{2}$ to bound the distance to net, Eq. (F.3) changes to be:
@@
where the first step follows from Lemma F.7, the second step follows from $\|x\|_{1}\leq\sqrt{d}\cdot\|x\|_{2}$ for $x\in\mathbb{R}^{d}$, and the last step follows from $\ell_{2}$ $\epsilon_{0}$-net.
Then, Eq. (F.3) changes to be
@@
For $\ell_{2}$ distance, we also have $|N|\leq(5R/\epsilon_{0})^{d}$ in Fact F.4.
We can choose $\epsilon_{0}=\Theta(1/n)$ to hide $nR_{w}\cdot\sqrt{d}\cdot\epsilon_{0}$ term in $B$ in Lemma F.8. Thus,
@@
Lemma F.11.
If we use $\ell_{1}$ $\epsilon_{0}$-net and use $\ell_{1}$ Lipschitz in Lemma F.8, we have copies of data structure
$l=O(d\log(nRd/p_{f}))$
@@
If we use $\ell_{1}$ to bound the distance to net, Eq. (F.3) changes to be:
@@
where the first step follows from Lemma F.7, and the last step follows from $\ell_{1}$ $\epsilon_{0}$-net.
Then, Eq. (F.3) changes to be
@@
We can choose $\epsilon_{0}=\Theta(1/n)$ to hide $nR_{w}\cdot\epsilon_{0}$ term in $B$ in Lemma F.8. Thus,
@@
  • $\ell_{\infty}$ $\epsilon_{0}$-net, Lemma F.9: we have $l=O(d\log(nR/p_{f}))$.
  • $\ell_{2}$ $\epsilon_{0}$-net, Lemma F.10: we have $l=O(d\log(nR/p_{f}))$.
  • $\ell_{1}$ $\epsilon_{0}$-net, Lemma F.11: we have $l=O(d\log(nRd/p_{f}))$.
@@
Appendix G Softmax Activation
In this section, we introduce how we extend previous $\ell_{1}$ distance results to the softmax activation function, which is the most widely used distance measure in attention mechanism based models. In order to have softmax activation, we need the distance measure function $\phi$ in Definition 1.1 to be the softmax function in Definition 4.1.
In Section G.1, we show how to extend to the softmax distance function in Lemma G.6. In Section G.2, we show how to adjust our algorithms. In Section G.3, we extend our algorithm to be robust to adaptive query. In Section G.4, we give the proof of our main result Theorem 5.1.
G.1 Exponential Inner Product
In this section, we show how we obtain the softmax distance using $\ell_{2}^{2}$ distance query. First, we provide some helpful results from [6].
Definition G.1 (Definition 3.1 in [6]).
@@ G.1 Exponential Inner Product
such that: There is a matrix $\widetilde{A}\in\mathbb{R}^{n\times n}$ that is an $(\epsilon,r)$-approximation (Definition G.1) of $A\in\mathbb{R}^{n\times n}$. Furthermore, the matrices $U_{1}$ and $U_{2}$ defining $\widetilde{A}$ can be computed in $O(n\cdot r)$ time.
Here we consider the vector version of Lemma G.2.
Definition G.3.
@@
Then, we have $P(x):[0,R]^{d}\to[0,\Gamma_{R,s}]^{r}$ where $P(\cdot)$ is polynomial kernel function defined in [6].
Remark G.4.
@@
Lemma G.5 (Polynomial approximation).
For any accuracy parameter $\epsilon_{s}\in(0,0.1)$, let $R\geq 1$, and let $P(x):[0,R]^{d}\to[0,\Gamma_{R,s}]^{r}$ be the $s$-th order polynomial kernel function defined in [6] where $r\leq{2s+2d\choose 2s}$ and $s=O(\max\{\frac{\log(1/\epsilon_{s})}{\log(\log(1/\epsilon_{s})/R)},R^{2}\})$. Then, for any $x,y\in[0,R]^{d}$, we have
@@
Let $n=1$. The proof follows from directly applying Lemma G.2. ∎
Using the results from [6] above, we can extend our results to softmax activation.
Lemma G.6 (Weighted softmax approximation, formal version of Lemma 6.7).
Let accuracy parameter be $\epsilon_{s}\in(0,0.1)$. Let $R\geq 1$. Let $r\leq{2s+2d\choose 2s}$ and $s=O(\max\{\frac{\log(1/\epsilon_{s})}{\log(\log(1/\epsilon_{s})/R)},R^{2}\})$. Let $P(x):[0,R]^{d}\to[0,\Gamma_{R,s}]^{r}$ be the $s$-th order polynomial kernel function defined in Lemma G.5. Then we can approximate exponential inner product using polynomial kernel function:
@@
From Lemma G.5, we can use polynomial kernel to approximate the softmax function:
@@
The proof of approximation error and time complexity of constructing $P(\cdot)$ follows from Lemma G.5.
@@
G.2 Algorithm Modifications
Based on Lemma G.6, we can now extend our DP algorithms to handle softmax activation. First, we need to construct $P(y)$ and $P(x_{i})$ for $i\in[n]$, each costing $O(r)$ time. Then, for the second term in Lemma G.6, i.e. $\frac{1}{2}\sum_{i\in[n]}w_{i}(\|P(x_{i})\|^{2}_{2}+\|P(y)\|^{2}_{2})$, we don’t need to add DP noises in it; instead, we calculate this term exactly, preprocess it, and store the results in the algorithm. For the first term, $-\frac{1}{2}\sum_{j\in[r]}\sum_{i\in[n]}w_{i}|P(x_{i})_{j}-P(y)_{j}|^{2}$, we can adjust our high dimensional DP distance query algorithm to solve it.

### v2

@@ G.2 High Dimension Single Data Structure
Theorem G.3 (DPTreeHighDim data structure).
@@
    • the process of output $z$ satisfies is $(\epsilon,\delta+\delta^{\prime})$-DP private, which computes $\sum_{i\in[n]}w_{i}\|y-x_{i}\|_{1}$,
@@
For the runtime analysis, since we loop data structure DPTreeDistance $d$ times, an additional $d$ factor will appear for both initialization and query time complexity. The DP is proved by Lemma G.1. The accuracy is proved by Lemma G.2. ∎
Appendix H Adaptive Query
In this section, we introduce how we can solve the adaptive query problem by our algorithm, using some tools from QRS+ [86]. Our idea is that, if we can prove that our algorithm can solve any query in the query space with certain error. Then, since adaptive query must lie in this space, we can handle adaptive query. In Section H.1, we show how we can boost the constant probability of our algorithm to high probability. In Section H.2, we show how we can apply the notion of $\epsilon_{0}$-net and bound all query points in net. In Section H.3, we show how we can bound all points in the query space by introducing an additive error. In Section H.4, we examine the effects of different norms on our adaptive query proof.
First, from Theorem G.3, given query $y\in[0,R]^{d},\alpha\in(0,1)$ we have DistanceQuery$(y,\alpha)$ that can solve $d$-dimension weighted $\ell_{1}$-distance problem with constant probability $0.99$. Now we show how to improve it to solve adaptive query problem.
H.1 Boost the Constant Probability to High Probability
We can repeat the data structure multiple times and take the median to boost the constant probability using Chernoff bound from Lemma B.2.
Lemma H.1 (Using Chernoff bound to boost the probability).
@@ H.1 Boost the Constant Probability to High Probability
Then for each fixed query point $y$, we can have the process of outputting the median of $l$ responses is $(\epsilon,\delta+\delta^{\prime})$-DP and the error is upper bounded by $\alpha A_{*}+B$ with probability $1-p_{f}$.
@@
By basic composition Fact B.6, we prove the DP. Similar to the proof of Theorem D.2, we prove the error by Chernoff bound (Lemma B.2). ∎
H.2 From Each Fixed Query Point to All On-net Points
@@
Definition H.2 ($\ell_{p}$ $\epsilon_{0}$-net, see Definition 4.2.1 in Ver [96]).
@@
Fact H.3 ($\ell_{\infty}$ $\epsilon_{0}$-net).
@@
Fact H.4 ($\ell_{2}$ $\epsilon_{0}$-net, see Lemma 5 in Woo [102]).
@@
Fact H.5 ($\ell_{1}$ $\epsilon_{0}$-net, see Theorem 2 in GS [41]).
@@
Lemma H.6 (From for each query point to for all points in net).
@@
Then with probability $1-p_{f}$, for all query points $y\in N$, we can have the process of outputting the median of $l$ responses is $(\epsilon,\delta+\delta^{\prime})$-DP and the error is upper bounded by $\alpha A_{*}+B$.
@@
By basic composition Fact B.6, we prove the DP. From Lemma H.1, we know for each $y\in N$, the error is upper bounded by $\alpha A_{*}+B$ with probability $1-p_{f}/|N|$.
@@
H.3 From Net Points to All Points
@@
Lemma H.7 (Lipschitz of query function).
@@
Lemma H.8 (From points in net to all points in query space).
@@
Then with probability $1-p_{f}$, for all query points $q\in{\mathcal{B}}$, there exists a point $y\in N$ which is the closest to $q$, we can have the process of outputting the median of $l$ responses is $(\epsilon,\delta+\delta^{\prime})$-DP and the error satisfy
@@
By basic composition Fact B.6, we prove the DP.
@@
From Lemma H.1, with $l=O(\log(|N|/p_{f}))$ we know
@@
where the first step follows from definition of $l$, the second step follows from Fact H.3.
We condition on event $E$ to be held. Then, by definition of $\ell_{\infty}$ $\epsilon_{0}$-net (see Definition H.2), for each $q\notin N$, there exists $y\in N$ such that
@@
where the first step follows from Lemma H.7, the second step follows from $\|x\|_{1}\leq d\|x\|_{\infty}$ for $x\in\mathbb{R}^{d}$, and the last step follows from Eq. (6).
@@
where the first step follows from triangular inequality, the second step follows from Eq. (H.3), the third step follows from Lemma H.6, and the last step follows from Eq. (H.3).
@@
H.4 Effect of Different Norms on the Result
In the above proof, we have two different measure spaces, i.e. $\ell_{\infty}$ distance of $\epsilon_{0}$-net (Definition H.2) and $\ell_{1}$ Lipschitz (Lemma H.7).
@@ H.4 Effect of Different Norms on the Result
For different norms, the only differences in the proofs will be Lipschitz smoothness in Eq. (H.3) and the cardinality of $\epsilon_{0}$-net, i.e. $|N|$ in Fact H.3.
Lemma H.9.
If we use $\ell_{\infty}$ $\epsilon_{0}$-net and use $\ell_{1}$ Lipschitz in Lemma H.8, we have copies of data structure $l=O(d\log(nR/p_{f}))$.
@@
If we use $\ell_{\infty}$ to bound the distance to net, Eq. (H.3) is:
@@
where the first step follows from Lemma H.7, the second step follows from $\|x\|_{1}\leq d\|x\|_{\infty}$ for $x\in\mathbb{R}^{d}$, and the last step follows from $\ell_{\infty}$ $\epsilon_{0}$-net.
Then, Eq. (H.3) is
@@
For $\ell_{\infty}$ distance, we have $|N|\leq(5R/\epsilon_{0})^{d}$ in Fact H.3.
We can choose $\epsilon_{0}=\Theta(1/n)$ to hide $nR_{w}\cdot d\cdot\epsilon_{0}$ term in $B$ in Lemma H.8. Thus,
@@
Lemma H.10.
If we use $\ell_{2}$ $\epsilon_{0}$-net and use $\ell_{1}$ Lipschitz in Lemma H.8, we have copies of data structure $l=O(d\log(nR/p_{f}))$.
@@
If we use $\ell_{2}$ to bound the distance to net, Eq. (H.3) changes to be:
@@
where the first step follows from Lemma H.7, the second step follows from $\|x\|_{1}\leq\sqrt{d}\cdot\|x\|_{2}$ for $x\in\mathbb{R}^{d}$, and the last step follows from $\ell_{2}$ $\epsilon_{0}$-net.
Then, Eq. (H.3) changes to be
@@
For $\ell_{2}$ distance, we also have $|N|\leq(5R/\epsilon_{0})^{d}$ in Fact H.4.
We can choose $\epsilon_{0}=\Theta(1/n)$ to hide $nR_{w}\cdot\sqrt{d}\cdot\epsilon_{0}$ term in $B$ in Lemma H.8. Thus,
@@
Lemma H.11.
If we use $\ell_{1}$ $\epsilon_{0}$-net and use $\ell_{1}$ Lipschitz in Lemma H.8, we have copies of data structure $l=O(d\log(ndR/p_{f}))$.
@@
If we use $\ell_{1}$ to bound the distance to net, Eq. (H.3) changes to be:
@@
where the first step follows from Lemma H.7, and the last step follows from $\ell_{1}$ $\epsilon_{0}$-net.
Then, Eq. (H.3) changes to be
@@
We can choose $\epsilon_{0}=\Theta(1/n)$ to hide $nR_{w}\cdot\epsilon_{0}$ term in $B$ in Lemma H.8. Thus,
@@
  • $\ell_{\infty}$ $\epsilon_{0}$-net, Lemma H.9: we have $l=O(d\log(nR/p_{f}))$.
  • $\ell_{2}$ $\epsilon_{0}$-net, Lemma H.10: we have $l=O(d\log(nR/p_{f}))$.
  • $\ell_{1}$ $\epsilon_{0}$-net, Lemma H.11: we have $l=O(d\log(nRd/p_{f}))$.
@@
Appendix I Softmax Activation
In this section, we introduce how we extend previous $\ell_{1}$ distance results to the Softmax activation function, which is the most widely used distance measure in attention mechanism based models.
In Section I.1, we show how to extend to the Softmax distance function in Lemma I.6. In Section I.2, we show how to adjust our algorithms. In Section I.3, we extend our algorithm to be robust to adaptive query. In Section I.4, we give the proof of our main result Theorem 3.1.
I.1 Exponential Inner Product
In this section, we show how we obtain the Softmax distance using $\ell_{2}^{2}$ distance query. First, we provide some helpful results from AS [6].
Definition I.1 (Definition 3.1 in AS [6]).
@@ I.1 Exponential Inner Product
such that: There is a matrix $\widetilde{A}\in\mathbb{R}^{n\times n}$ that is an $(\epsilon,r)$-approximation (Definition I.1) of $A\in\mathbb{R}^{n\times n}$. Furthermore, the matrices $U_{1}$ and $U_{2}$ defining $\widetilde{A}$ can be computed in $O(n\cdot r)$ time.
Here we consider the vector version of Lemma I.2.
Definition I.3.
@@
Then, we have $P(x):[0,R]^{d}\to[0,\Gamma_{R,s}]^{r}$ where $P(\cdot)$ is polynomial kernel function defined in AS [6].
Remark I.4.
@@
We use the polynomial approximation method, which has been applied to accelerate Transformer model extensively [6, 7, 8, 72, 69].
Lemma I.5 (Polynomial approximation).
For any accuracy parameter $\epsilon_{s}\in(0,0.1)$, let $R\geq 1$, and let $P(x):[0,R]^{d}\to[0,\Gamma_{R,s}]^{r}$ be the $s$-th order polynomial kernel function defined in AS [6] where $r\leq{2s+2d\choose 2s}$ and $s=O(\max\{\frac{\log(1/\epsilon_{s})}{\log(\log(1/\epsilon_{s})/R)},R^{2}\})$. Then, for any $x,y\in[0,R]^{d}$, we have
@@
Let $n=1$. The proof follows from directly applying Lemma I.2. ∎
Using the results from AS [6] above, we can extend our results to Softmax activation.
Lemma I.6 (Weighted Softmax approximation, formal version of Lemma C.7).
Let accuracy parameter be $\epsilon_{s}\in(0,0.1)$. Let $R\geq 1$. Let $r\leq{2s+2d\choose 2s}$ and $s=O(\max\{\frac{\log(1/\epsilon_{s})}{\log(\log(1/\epsilon_{s})/R)},R^{2}\})$. Let $P(x):[0,R]^{d}\to[0,\Gamma_{R,s}]^{r}$ be the $s$-th order polynomial kernel function defined in Lemma I.5. Then we can approximate exponential inner product using polynomial kernel function:
@@
From Lemma I.5, we can use polynomial kernel to approximate the Softmax function:
@@
The proof of approximation error and time complexity of constructing $P(\cdot)$ follows from Lemma I.5.
@@
I.2 Algorithm Modifications
Based on Lemma I.6, we can now extend our DP algorithms to handle Softmax activation. First, we need to construct $P(y)$ and $P(x_{i})$ for $i\in[n]$, each costing $O(r)$ time. Then, for the second term in Lemma I.6, i.e. $\frac{1}{2}\sum_{i\in[n]}w_{i}(\|P(x_{i})\|^{2}_{2}+\|P(y)\|^{2}_{2})$, we don’t need to add DP noises in it; instead, we calculate this term exactly, preprocess it, and store the results in the algorithm. For the first term, $-\frac{1}{2}\sum_{j\in[r]}\sum_{i\in[n]}w_{i}|P(x_{i})_{j}-P(y)_{j}|^{2}$, we can adjust our high dimensional DP distance query algorithm to solve it.